What is the axis of symmetry for the graph of y=2x2-4x+2


1
Expert's answer
2022-03-14T19:05:59-0400

Since

"y=2x^2-4x+2\n=2(x^2-2x+1)\n=2(x-1)^2,"

we conclude that the axis of symmetry is "x=1."